OpenStudy (anonymous):

find the slope of a line perpendicular to x-y=16

OpenStudy (anonymous):

-1

OpenStudy (anonymous):

If two lines are perpendicular they have the following property:\[m _{1}m _{2}=-1\]Where m1 and m2 are the slopes of the two lines. Your line has a slope of 1 thus we solve for m2 to get the slope of a line perpendicular to your line:\[m _{2}=-1/m _{1}=-1/1=-1\]